OpenStudy (anonymous):

what is the lcd of 7/3x+2/x+x/x-1 would it be 3x-1?

hero (hero):

Use the fraction function in equation editor to post your question.

OpenStudy (anonymous):

\[\frac{ 7 }{ 3x } + \frac{ 2 }{ x } + \frac{ x }{ x-1 } ?\]

hero (hero):

I suppose that's what it is. I couldn't really tell at first

hero (hero):

But the LCD is not 3x-1

hero (hero):

It's more like \(3x(x-1)x\)

OpenStudy (anonymous):

oh okay thank you